OpenStudy (tiffanymak1996):

find limit of {pi^n / 5^n} n=1 , +infinite

Parth (parthkohli):

Actually, this is a continuous function.

Parth (parthkohli):

\[\lim_{x \to + \infty} {\pi^n \over 5^{n}} = {} \]

Parth (parthkohli):

Observe the graph, and tell me what \(f(x)\) approaches when \(x\) approaches infinity.

Parth (parthkohli):

Wait, what is that \(n =1,+\infty\)?

Parth (parthkohli):

\[\lim_{n \to 0 } ({\pi \over 5})^n \]

OpenStudy (tiffanymak1996):

it's sequences

Parth (parthkohli):

It's a function, tiffany.

Parth (parthkohli):

Where does the function approach?

Parth (parthkohli):

Maybe it's a sequence. Can you just draw it out?

OpenStudy (tiffanymak1996):